Redesigned Computers has 6.5 percent coupon bonds outstanding with a maturity date of 8 years from now. The yield to maturity is 16 percent and the face value is $1,000. Interest is paid semiannually. What is the market price for these bonds?
Answer #1

Number of periods = 8 * 2 = 16

Rate = 16% / 2 = 8%

Coupon = (0.065 * 1000) / 2 = 32.5

Market price = Coupon *[1- 1 / (1 + r)n] / r + FV/ (1 + r)n

Market price = 32.5 * [1- 1 / (1 + 0.08)16] / 0.08 + 1000 / (1 + 0.08)16

Market price = 32.5 * [1 - 0.29189] / 0.08 + 291.89047

Market price = 32.5 * 8.85137 + 291.89047

Market price = $579.56